OpenStudy (anonymous):

How many years will it take for a $50 investment to grow to $5000 if it is compounded continuously at a rate of 5%? Round to two decimal places. Do not include units in your answer A(t) = P*e^(r*t)

OpenStudy (amistre64):

5000 = 50(e^(.05t)) 500/50 = e^(.05t) ln(100) ------ = t .05

OpenStudy (anonymous):

so whats the answer

OpenStudy (amistre64):

this is a study site, not a free answering service ....

OpenStudy (anonymous):

ok?